OpenStudy (anonymous):

Fool's problem of the day, A function \( f \) is defined such that \( f(2) = 60 \) and \( \sum \limits_{i=1} ^n (-1)^i f(i) = nf(n) \forall n > 1 \text{ and } n \in \mathbb{N} \). Can you find \( f(257) \) ? [Solved by @dumbcow] Good luck!

OpenStudy (lgbasallote):

i just came here to give the medal. byeeee

OpenStudy (dumbcow):

seems like as n increases f(n) decreases toward 0

OpenStudy (perl):

-1 f(1) = 1 f(1) -> f(1) = 0 -1f(1) + f(2) = 2 f(2) -1 f(1) + 60 = 2 f(2)

OpenStudy (apoorvk):

yeha increase and decrease.. but i can t seem to find a trend.. calculating till 57 will take a lot of time... :/ may be we need to go reverse.

OpenStudy (dumbcow):

f(11) = f(12) = 10 thats as far as i've gotten

OpenStudy (apoorvk):

i found all values till f(9). but we slog more? am sure there's more to it.

OpenStudy (perl):

did you get zero for f(1)

OpenStudy (apoorvk):

and f(3) = f(4) = 30 no f(1) i got -60 if i am right at all

OpenStudy (dumbcow):

no i got f(1) = -60 --> -(-60) +60 = 2*60

OpenStudy (apoorvk):

i even have the general equation of every term implicit form.. but what then?

OpenStudy (anonymous):

@dumbcow: The original problem ask for f(11) only, but i derived the a closed form for f(n) and revised the constraints ;)

OpenStudy (anonymous):

@apoorvk: You would get a recurrence relation then you can use various method to get the closed form in terms of f(2). Okay no more hints.

OpenStudy (dumbcow):

i see, yeah thats the hard part getting the closed form for f(n)

OpenStudy (perl):

something seems odd, if you plug in n=1 , or you cant sum from i = 1 to n=1 ?

OpenStudy (anonymous):

perl, the function f(n) is defined only for for n>1, my apologies.

OpenStudy (dumbcow):

i noticed something else, the sum seems to stay constant at 120 when n is odd so then it follows 257*f(257) = 120-f(257) f(257) = 120/258 = .465 is that right?

OpenStudy (apoorvk):

\[(\sum_{i=1}^{n-1} (-1)^i f(i)) / (n +1) = f(n)\] .. when n is odd \[(\sum_{i=1}^{n-1} (-1)^i f(i)) / (n -1) = f(n)\] .. when n is even

OpenStudy (anonymous):

Congratz @dumbcow that's the right answer.

OpenStudy (apoorvk):

yeah sorry when n is 'odd' and so sum of (n-1) terms is 120. so according to the equation i posted, f(257), which is odd, is 120/(257+1) =120/258 =0.46512 Dumbcow did it first!!!!!!!!

OpenStudy (anonymous):

That's it, just one problem? :/